# On Bureau’s classification of quadratic differential equations in two variables free of movable critical points

### v2

@@ 2.2. Canonical equations and Malmquist’s conditions
and for $z$ defined by $x=(z-u)y$, the system becomes
@@ 3. The classification
We follow Bureau’s steps for the greater part of the classification. We start with an equation of the form
@@
free of movable critical points, with $P$ and $Q$ quadratic homogeneous polynomials in $y$ and $z$ (at least one of them non-trivial), whose coefficients are holomorphic functions of the independent variable, defined in a domain $U\subset\mathbf{C}$. By the results in Section 2.1, for all $t$ in $U$, the linear equivalence class of the quadratic part of (6) is represented by one of the autonomous systems of Table 2, and is thus, with the exception of a discrete set of values of $t$, locally constant (this is the first source of exceptional values of $t$ in Theorem 1). By restricting $t$ to a neighborhood of such a value, up to a change of coordinates that is linear in $y$ and $z$, and that preserves the independent variable, we may suppose that the quadratic part in (6) is independent of $t$, and that it is one of the quadratic systems of Table 2.
@@
where the parameters $\lambda$, $\mu$, $\nu$, $h$ and $k$ are functions of $t$, chosen to solve a particular system of algebro-differential equations in a neighborhood of some $t_{0}\in U$ (the need, in Theorem 1, to restrict the domains is also a consequence of the local character of this analysis).
@@
the system will have poles along the line at infinity for all values of $t$, with singularities at the directions of the radial orbits of the quadratic part. At a singularity associated to a radial orbit with positive index $n$, we will obtain a canonical equation (5) with index $n$, and the results in Section 2.2 will impose a necessary condition for the absence of movable critical points, of algebro-differential nature, which may be explicitly calculated after performing $n-1$ blow-ups. We will meet some changes of coordinates that may be framed within this setting; for instance, the coordinates $(1/y,z)$ amount to considering, first, a change of affine chart in $\mathbf{P}^{2}$, $(u,v)=(1/y,z/y)$, followed by a blow-up, $(u,v)\mapsto(u,v/u)=(1/y,z)$.
@@ 3.1. System I
$(y(t),z(t))=\left(\frac{\alpha y_{0}+\beta z_{0}+\gamma}{\eta y_{0}+\theta z_{0}+\iota},\frac{\delta y_{0}+\epsilon z_{0}+\zeta}{\eta y_{0}+\theta z_{0}+\iota}\right),$
with $\alpha$, $\beta$, etc. functions of $t$, and these equations are hence naturally free of movable critical points in all the domain where they are defined. A linear transformation redresses $L(y,z,t)$ into $-y$, and, after this, a suitable transformation $\text{T}_{3}$ brings the system into
@@
(a form simpler than Bureau’s, who uses the more restrictive transformation $\text{T}_{2}$). Explicit expressions for its integration follow from the fact that $(1/z)^{\prime\prime}=B$.
@@ 3.2. System II
For $y=1/u$ and $z=v/u$ we have the canonical equation of index $1$
@@
From Proposition 6, $C=0$ is a necessary condition for the absence of movable critical points. We will henceforth assume it. At the point $(u,v)=(0,1)$, the system is a canonical equation of index $1$, from which we obtain the necessary condition $A=0$. With it, $u=y-z$ solves the Riccati equation with holomorphic coefficients $u^{\prime}=u^{2}+a$, which is free of movable critical points. Thus, the system solved by $u$ and $z$ (and hence the one solved by $y$ and $z$) is free of movable critical points.
@@ 3.3. System V
($f^{\prime}$ denotes the derivative of $f$ with respect to $t$.) This shows that, in (8), the condition $B=0$ is, with respect to affine changes of coordinates, an intrinsic one. We leave out of our analysis the values of $t$ where $B$ has an isolated zero, and restrict to domains where either $B$ vanishes identically, or where it does nor vanish at all (this is another source of special values of the independent variable in the statement of Theorem 1; conditions of the sort will be implicitly assumed throughout the rest of the article).
@@
Bureau obtains the necessary condition $f^{\prime\prime}=0$ through a canonical equation obtained from Eq. (9). This works both ways: Bureau’s analysis implies that, in Eq. (10), $f^{\prime\prime}=0$ is a necessary condition for the absence of movable critical points. We will reobtain this condition in sections 3.7.2.1, 3.7.2.3 and 3.11.
@@ 3.4. System VI
by a transformation of type $\text{T}_{2}$. In the coordinates $u=-1/z$ and $v=y/z$, we find the canonical equation of index $1$,
@@ 3.5. System VII
In the coordinates $(u,v)$ defined by $y=-1/u$, $z=(v+\frac{1}{2})/u^{2}$, we find the canonical system of index $2$
@@ 3.6. System VIII
Studied by Bureau in [Bur81b, §17]. It has a parameter $n\in\mathbf{Z}$, $n\geq 1$. By means of a transformation $\text{T}_{2}$, it may be brought into Bureau’s form
@@
In the coordinates $(u,v)$ defined by $y=1/u$, $z=(v+1)/u$, we have the canonical system of index $1$
@@
A first necessary condition for the absence of movable critical points is thus $A=0$, which we assume from now on. For $y=-1/u$, we have the canonical equation of index $n$
@@
for $\text{for }n=1,\quad$ $b=0;$
for $\text{for }n=2,\quad$ $b^{\prime}=0;$
for $\text{for }n=3,$ $b^{\prime\prime}+ab-3b^{2}=0;$
for $\text{for }n=4,$ $b^{\prime\prime\prime}+4(a-4b)b^{\prime}+2ba^{\prime}=0;$
@@
In order to integrate the system (12) with $A=0$, Bureau transforms it into a particular case of the second-order equation in one variable studied by Gambier in [Gam10, Ch. VII, Eq. 5, p. 53], which he revisits in [Bur64, §49] (when $a\neq 0$, he passes through the second-order equation satisfied by $1/y$; when $n\neq 2$ and $b\neq 0$, through the one satisfied by $z$; he deals separately with the case $n=2$). Gambier’s equation has a rather convoluted scheme of integration. In it, the necessary and sufficient conditions for the absence of movable critical points are easy to establish for the equations corresponding to low values of $n$; for higher values, either the absence of movable critical points can be established directly, through the vanishing of a particular explicit expression, or the equation may be transformed into one of the same family, in which $n$ is replaced by $n-2$. We refer to Gambier’s work for details.
The eventual movable critical points of the system are easy to locate. For $z$ and $w=y-z$, we have the system
@@
The system has poles along $u=0$, and nowhere else. In the chart (14), the equation has a singularity with index $n$ along $p:(u,z)=(0,0)$, and, in the chart (15), a singularity with index $-n$ at $q:(u,v)=(0,0)$.
Elementary transformations play an important role in the birational theory both of ruled surfaces and of the Riccati equation (see [BHPV04, Ch. 6, Sect. 7] and [Bru15, Ch. 4, §1]; elementary transformations are called “flipping the fiber” in the latter). We will perform parametric (non-autonomous) versions of such transformations.
Start with the rational fibration $\Pi:\mathbf{P}^{1}\times\mathbf{P}^{1}\times U\to\mathbf{P}^{1}\times U$, $(1:u,z:1,t)\mapsto(1:u,t)$. Let $F$ be the fiber of $\Pi$ above $\{u=0\}\times U$. For the intersection with a slice $\Sigma:t=t_{0}$, $F\cap\Sigma$ is, within $\Sigma$, a rational curve of vanishing self-intersection. The blowing up of $p$ produces a surface $E$, and transforms $F$ into a surface $F^{\prime}$. Within the transform of $\Sigma$, both $E\cap\Sigma$ and $F^{\prime}\cap\Sigma$ are rational curves of self-intersection $-1$. By the parametric version of Grauert’s criterion due to Nakano and Fujiki [Nak71, FN72], $F^{\prime}$ may be contracted along the fibers of $t$, producing a curve $q^{\prime}$ transverse to $t$. This transforms $E$ into a surface $E^{\prime}$, and, within the new transform of $\Sigma$, $E^{\prime}\cap\Sigma$ is a curve of vanishing self-intersection. The total space $\mathbf{P}^{1}\times\mathbf{P}^{1}\times U$ has been modified into a threefold $M$, and the transform of $\Pi$ subsists as a locally trivial rational bundle $\Pi^{\prime}:M\to\mathbf{P}^{1}\times U$. The fiber $F$ has been replaced by $E^{\prime}$, but all the other fibers have gone by unchanged. The transformed equation has poles along $E^{\prime}$. There is a curve $p^{\prime}$ on $E^{\prime}$, along which we have a canonical equation with index $n-1$; at the curve $q^{\prime}$, we find a canonical equation with index $1-n$. This procedure can be iterated: when applied to a point of index $1$, the conditions of Proposition 6 are equivalent to the fact that the poles in the equation disappear after blowing up the point, thus removing all poles from the equation.
After iterating the procedure, $\mathbf{P}^{1}\times\mathbf{P}^{1}\times U$ has been birationally transformed into a threefold where the equation extends holomorphically (where it no longer has poles), such that the fibers of the projection onto $t$ compactify the original $\mathbf{C}^{2}\times\{t\}$. This implies that the equation is free of movable critical points (see [Oka79, §1], where Okamoto resorts to Ehresmann’s theorem, or Painlevé’s 17th of his Leçons de Stockholm [Pai97]).
@@ 3.7. System IX (plus systems III and IV)
with $n\in\mathbf{Z}$, $n\neq 0$. System III corresponds to $n=-1$, system IV to $n=1$, and system X to the remaining cases. The latter is split in two: system IX.A corresponds to $n<-1$, and system IX.B to $n>1$. Warning: in system IX.A, Bureau’s normalization differs from the one here used by the sign of $n$.
Let us be overly explicit in the calculations in this part, in which Bureau’s analysis presents some oversights. Under a transformation of type $\text{T}_{1}$, in the form
@@ 3.7.2. Case $B\neq 0$, $n\neq-3$
$-\frac{v^{\prime\prime}}{(v^{\prime})^{2}}=-\frac{3v-n}{v(2v-n-1)},$ (20)
discussed in Section 2.1. We must impose the condition $n\neq-1$ in order for the pole at $v=0$ to be simple. The expression (20) has a simple pole at
$v_{1}=\frac{1}{2}(n+1),$
@@
If $v$ is univalent, the latter is either equal to $-1$ or of the form $11/k-$ 1 for $k\in\mathbf{Z}$, this is, $2(n+1)/(n-1)=2+4/(n-1)$ is an integer (the index). The admissible cases are $n\in\{2,3,5\}$. The indices of the poles of (20) appear in Table 3.
$n$ 0 $v_{1}$infty
$2$ $3$ $6$ $2$
$3$ $4$ $4$ $2$
$5$ $6$ $3$ $2$
@@
We have two points along $u=0$: $v=0$ and $v=v_{1}$. At each one of them we have a canonical equation whose index appears in Table 3. This will give, in each of the admissible cases, two necessary conditions for the absence of movable critical points.
3.7.2.1 Subcase $n=2$
The necessary condition coming from (21) along $(u,v)=(0,0)$ is $b=0$, which we assume from now on. Set $a=12q$. For $u=\frac{1}{6}z+q$ and $f=q^{\prime\prime}-6q^{2}$, we have that $u^{\prime\prime}=6u^{2}+f$, the equation appearing in Section 3.3 as Eq. (10), related to the first Painlevé equation. As we there discussed, the necessary and sufficient condition for the absence of movable critical points is $f^{\prime\prime}=0$; this is also the necessary condition arising from Eq. (21) at $(u,v)=(0,v_{1})$.
The equation is birationally equivalent to the natural first-order system in one variable associated to the first Painlevé equation: $u$ and $u^{\prime}$ may be rationally expressed in terms of $y$ and $z$; on their turn, from (18), $y$ and $z$ may be rationally expressed in terms of $u$ and $u^{\prime}$.
3.7.2.2 Subcase $n=3$
We have, for $f=-2a$ and $g=b+a^{\prime}$,
@@ 3.7.2.2 Subcase $n=3$
We find in [Inc44, §14.315, p. 333] that the conditions for this equation to be free of movable critical points are $f^{\prime\prime}=0$ and $g^{\prime}=0$ ($a^{\prime\prime}=0$ and $b^{\prime}=0$). These are exactly the two conditions arising from equation (21) (they will be also discussed in Section 3.10). Under these, (22) is the general form of the equation leading to the Painlevé II transcendent [Pai02, §11]. If $f^{\prime}\neq 0$, up to a translation the independent variable and suitable scalings, equation (22) gives one of the standard forms of the Painlevé II equation
@@
When $f^{\prime}=0$, (22) is integrated by elliptic functions and their rational degenerations: if $f$ and $g$ are, respectively, the constants $f_{0}$ and $g_{0}$, we have the first integral $(y^{\prime})^{2}=y^{4}+f_{0}y^{2}+2g_{0}y+K$ (see also [Inc44, §14.315, p. 333]). When substituting for $y^{\prime}$ from (18), we obtain that
$2y^{2}z-z^{2}+2by-2az-a^{2}$
is a first integral for (18). Its generic level surfaces are elliptic fibrations over the domain of the independent variable.
The system is polynomially equivalent to the first-order system in two variables associated to the general Painlevé II equation: from (18), $y$ and $z$ may be polynomially expressed in terms of $y$ and $y^{\prime}$, and vice versa.
3.7.2.3 Subcase $n=5$
The necessary condition arising from Eq. (21) at $(u,v)=(0,v_{1})$ is $b=-3a^{\prime}$. Set $a=3q$ and $b=-9q^{\prime}$. For $u=-2y$, we have
$u^{\prime\prime}=-uu^{\prime}+u^{3}-12qu+12q^{\prime}.$ (23)
This equation is integrated in [Pai02, p. 23, Eq. 6]: if $w$ is a solution to
$w^{\prime\prime}-6w^{2}=q^{\prime\prime}-6q^{2}$ (24)
different from $q$,
$u=\frac{w^{\prime}-q^{\prime}}{w-q}$ (25)
is a solution to (23). When considered as first-order equations in two variables, equations (23) and (24) are birationally equivalent, since, from (25), $w$ and $w^{\prime}$ can be expressed rationally in terms of $u$ and $u^{\prime}$:
$\begin{split}w&=\textstyle{\frac{1}{6}}u^{2}-q+\textstyle{\frac{1}{6}}u^{\prime},\\ w^{\prime}&=\textstyle{\frac{1}{6}}u^{3}-2uq+\textstyle{\frac{1}{6}}uu^{\prime}+q^{\prime},\end{split}$
@@ 3.7.2.3 Subcase $n=5$
in which (24) becomes the first Painlevé equation or one of its degenerations. Thus, condition (26), is a necessary and sufficient condition for (23) to be free of movable critical points as well (compare with [Inc44, §14.314, p. 331–333]). Condition (26) is also the one arising from the canonical equation (21) at $(u,v)=(0,0)$.
3.7.3. Case $B\neq 0$, $n=-3$

### v3

@@ 2.2. Canonical equations and Malmquist’s conditions
and for $z$ defined by x=(zu)yx=(z-u)y, the system becomes
@@ 3. The classification
We follow Bureau’s steps for the greater part of the classification. We start with a system of equations of the form
@@
with $P$ and $Q$ quadratic homogeneous polynomials in $y$ and $z$ (at least one of them non-trivial), whose coefficients, along with $A$, $B$, $C$, $D$, $a$ and $b$, are holomorphic functions of the independent variable defined in a domain $U\subset\mathbf{C}$, and we suppose the system to be free of movable critical points. By the results in Section 2.1, for all $t$ in $U$, the linear equivalence class of the quadratic part of (10) is represented by one of the autonomous systems of Table 3, and is thus, with the exception of a discrete set of values of $t$, locally constant (this is the first source of exceptional values of $t$ in Theorem 1). By restricting $t$ to a neighborhood of such a value, up to a change of coordinates that is linear in $y$ and $z$, and that preserves the independent variable, we may suppose that the quadratic part in (10) is independent of $t$, and that it is one of the quadratic systems of Table 3.
@@
where $\lambda$, $\mu$, $\nu$, $h$ and $k$ are functions of $t$, chosen to solve a particular system of algebro-differential equations in a neighborhood of some $t_{0}\in U$ (the need, in Theorem 1, to restrict the domains, is also a consequence of the local character of this analysis).
@@
the system will have poles along the line at infinity for all values of $t$, with singularities at the directions of the radial orbits of the quadratic part. At a singularity associated to a radial orbit with positive index $n$, we will obtain a canonical equation (9) with index $n$, and the results in Section 2.2 will impose a necessary condition for the absence of movable critical points, of algebro-differential nature, which may be explicitly calculated after performing n1n-1 blow-ups. We will meet some changes of coordinates that may be framed within this setting; for instance, the coordinates $(1/y,z)$ amount to considering, first, a change of affine chart in $\mathbf{P}^{2}$, $(u,v)=(1/y,z/y)$, followed by a blow-up, $(u,v)\mapsto(u,v/u)=(1/y,z)$.
@@ 3.1. System I
$(y,z)=\left(\frac{\alpha y_{0}+\beta z_{0}+\gamma}{\eta y_{0}+\theta z_{0}+\iota},\frac{\delta y_{0}+\epsilon z_{0}+\zeta}{\eta y_{0}+\theta z_{0}+\iota}\right),$
with $\alpha$, $\beta$, etc. functions of $t$, and these equations are hence naturally free of movable critical points throughout the domain where they are defined. A linear transformation redresses $L(y,z,t)$ into y-y, and, after this, a suitable transformation $\text{T}_{3}$ brings the system into
@@
(a form simpler than Bureau’s, who uses the more restrictive transformation $\text{T}_{2}$). An explicit expression for its integration follows from the fact that $(1/z)^{\prime\prime}=B$.
@@ 3.2. System II
For $y=1/u$ and $z=v/u$ we have the canonical equation of index 11
@@
From Proposition 6, $C=0$ is a necessary condition for the absence of movable critical points. We will henceforth assume it. At the point $(u,v)=(0,1)$, the system is a canonical equation of index $1$, from which we obtain the necessary condition $A=0$. With it, u=yzu=y-z solves the Riccati equation with holomorphic coefficients $u^{\prime}=u^{2}+a$, which is free of movable critical points. Thus, the system solved by $u$ and $z$ (and hence the one solved by $y$ and $z$) is free of movable critical points.
@@ 3.3. System V
($f^{\prime}$ denotes the derivative of $f$ with respect to $t$.) This shows that, in (12), the condition $B=0$ is, with respect to affine changes of coordinates, an intrinsic one. We leave out of our analysis the values of $t$ where $B$ has an isolated zero, and restrict to domains where either $B$ vanishes identically, or where it does not vanish at all (this is another source of special values of the independent variable in the statement of Theorem 1; conditions of the sort will be implicitly assumed throughout the rest of the article).
@@
Bureau obtains the necessary condition $f^{\prime\prime}=0$ through a canonical equation obtained from Eq. (13). This works both ways: Bureau’s analysis implies that, in Eq. (14), $f^{\prime\prime}=0$ is a necessary condition for the absence of movable critical points. We will reobtain this condition in sections 3.7 and 3.11.
@@ 3.4. System VI
by a transformation of type $\text{T}_{2}$. In the coordinates u=1/zu=-1/z and $v=y/z$, we find the canonical equation of index $1$,
@@ 3.5. System VII
In the coordinates $(u,v)$ defined by y=1/uy=-1/u, $z=(v+\frac{1}{2})/u^{2}$, we find the canonical system of index 22
@@ 3.6. System VIII
Studied by Bureau in [Bur81b, §17]. It has a parameter $n\in\mathbf{Z}$, $n\geq 1$. By means of a transformation $\text{T}_{2}$, the system may be brought into Bureau’s form
@@
In the coordinates $(u,v)$ defined by $y=1/u$, $z=(v+1)/u$, we have the canonical system of index 11
@@
A first necessary condition for the absence of movable critical points is thus $A=0$, which we assume from now on. For y=1/uy=-1/u, we have the canonical equation of index $n$
@@
$\text{for }n=1,\quad$ $b=0;$
$\text{for }n=2,\quad$ $b^{\prime}=0;$
$\text{for }n=3,$ $b^{\prime\prime}+ab-3b^{2}=0;$
$\text{for }n=4,$ $b^{\prime\prime\prime}+4(a-4b)b^{\prime}+2ba^{\prime}=0;$
@@
In order to integrate the system (16) with $A=0$, Bureau transforms it into a particular case of the second-order equation in one variable studied by Gambier in [Gam10, Ch. VII, Eq. 5, p. 53], which he revisits in [Bur64, §49] (when $a\neq 0$, he passes through the second-order equation satisfied by $1/y$; when $n\neq 2$ and $b\neq 0$, through the one satisfied by $z$; he deals separately with the case $n=2$). Gambier’s equation has a rather convoluted scheme of integration. In it, the necessary and sufficient conditions for the absence of movable critical points are easy to establish for the equations corresponding to low values of $n$; for higher values, either the absence of movable critical points can be established directly, through the vanishing of a particular explicit expression, or the equation may be transformed into one of the same family, in which $n$ is replaced by n2n-2. We refer to Gambier’s work for details.
The eventual movable critical points of the system are easy to locate. For $z$ and w=yzw=y-z, we have the system
@@
The system has poles along $u=0$, and nowhere else. In the chart (18), the equation has a singularity with index $n$ along $p:(u,z)=(0,0)$, and, in the chart (19), a singularity with index n-n at $q:(u,v)=(0,0)$.
Elementary transformations play an important role in the birational theory both of ruled surfaces and of the Riccati equation (see [BHPV04, Ch. 6, Sect. 7] and [Bru15, Ch. 4, §1]; they are called “flips of the fiber” in the latter). We will perform parametric (non-autonomous) versions of such transformations.
Start with the rational fibration $\Pi:\mathbf{P}^{1}\times\mathbf{P}^{1}\times U\to\mathbf{P}^{1}\times U$, $(1:u,z:1,t)\mapsto(1:u,t)$. Let $F$ be the fiber of $\Pi$ above $\{u=0\}\times U$. For the intersection with a slice $\Sigma:t=t_{0}$, $F\cap\Sigma$ is, within $\Sigma$, a rational curve of vanishing self-intersection. The blowing up of $p$ produces a surface $E$, and transforms $F$ into a surface $F^{\prime}$. Within the transform of $\Sigma$, both $E\cap\Sigma$ and $F^{\prime}\cap\Sigma$ are rational curves of self-intersection 1-1. By the parametric version of Grauert’s criterion due to Nakano and Fujiki [Nak71, FN72], $F^{\prime}$ may be contracted along the fibers of $t$, producing a curve $q^{\prime}$ transverse to $t$. This transforms $E$ into a surface $E^{\prime}$, and, within the new transform of $\Sigma$, $E^{\prime}\cap\Sigma$ is a curve of vanishing self-intersection. The total space $\mathbf{P}^{1}\times\mathbf{P}^{1}\times U$ has been modified into a threefold $M$, and the transform of $\Pi$ subsists as a locally trivial rational bundle $\Pi^{\prime}:M\to\mathbf{P}^{1}\times U$. The fiber $F$ has been replaced by $E^{\prime}$, but all the other fibers have gone by unchanged. The transformed equation has poles along $E^{\prime}$. There is a curve $p^{\prime}$ on $E^{\prime}$, along which we have a canonical equation with index n1n-1; at the curve $q^{\prime}$, we find a canonical equation with index 1n1-n. This procedure can be iterated: when applied to a point of index $1$, the conditions of Proposition 6 are equivalent to the fact that the poles in the equation disappear after blowing up the point, thus removing all poles from the equation.
After iterating the procedure, $\mathbf{P}^{1}\times\mathbf{P}^{1}\times U$ has been birationally transformed into a threefold where the equation extends holomorphically (where it no longer has poles), such that the fibers of the projection onto $t$ compactify the original fiber $\mathbf{C}^{2}\times\{t\}$. This implies that the equation is free of movable critical points (see [Oka79, §1], where Okamoto resorts to Ehresmann’s theorem, or Painlevé’s 17th of his Leçons de Stockholm [Pai97]).
@@ 3.7. System IX (plus systems III and IV)
with $n\in\mathbf{Z}$, $n\neq 0$. System III corresponds to n=1n=-1, system IV to $n=1$, and system X to the remaining cases. The latter is split in two: system IX.A corresponds to n<1n<-1, and system IX.B to $n>1$. Warning: in system IX.A, Bureau’s normalization differs from the one here used by the sign of $n$.
Let us be overly explicit in the calculations in this part, in which Bureau’s analysis presents some oversights. Under a transformation of type $\text{T}_{1}$, of the form
@@ 3.7.2. Case $B\neq 0$, $n\neq-3$
$\frac{v^{\prime\prime}}{(v^{\prime})^{2}}=-\frac{3v-n}{v(2v-n-1)},$ (24)
discussed in Section 2.1. We must impose the condition $n\neq-1$ in order for the pole at $v=0$ to be simple. The expression (24) has a simple pole at $v=v_{1}$ for
$v_{1}={\textstyle\frac{1}{2}}(n+1),$
@@
If $v$ is univalent, the latter is either equal to 1-1 or has the form 1/k11/k-1 for $k\in\mathbf{Z}$, this is, 2(n+1)/(n1)=2+4/(n1)2(n+1)/(n-1)=2+4/(n-1) is an integer (the index). The admissible cases are $n\in\{2,3,5\}$. The indices of the poles of (24) appear in Table 4.
$n$ 0 $v_{1}$ $\infty$
22 33 66 22
33 44 44 22
55 66 33 22
@@
We have two singular points along $u=0$: $v=0$ and $v=v_{1}$. At each one of them, we have a canonical equation whose index appears in Table 4. This will give, in each of the admissible cases, two necessary conditions for the absence of movable critical points.
Subcase $n=2$
The necessary condition coming from (25) along $(u,v)=(0,0)$ is $b=0$, which we assume from now on. Set $a=12q$. For $u=\frac{1}{6}z+q$, we obtain $u^{\prime\prime}-6u^{2}=q^{\prime\prime}-6q^{2}$, the equation appearing in Section 3.3 as Eq. (14) for $f=q^{\prime\prime}-6q^{2}$, related to the first Painlevé equation. As we there discussed, the necessary and sufficient condition for the absence of movable critical points is $f^{\prime\prime}=0$; this is also the necessary condition arising from Eq. (25) at $(u,v)=(0,v_{1})$.
The system is birationally equivalent to the first-order system in two variables associated to the first Painlevé equation (cf. Proposition 2).
Subcase $n=3$
We have, for f=2af=-2a and $g=b+a^{\prime}$,
@@
We find in [Inc44, §14.315, p. 333] that the conditions for this equation to be free of movable critical points are $f^{\prime\prime}=0$ and $g^{\prime}=0$ ($a^{\prime\prime}=0$ and $b^{\prime}=0$). These are exactly the two conditions arising from Eq. (25) (they will also be discussed in Section 3.10). Under these, (26) is the general form of the equation leading to the Painlevé II transcendent [Pai02, §11]. If $f^{\prime}\neq 0$, up to a translation the independent variable and suitable scalings, Eq. (26) gives one of the standard forms of the Painlevé II equation,
@@
When $f^{\prime}=0$, (26) is integrated by elliptic functions and their rational degenerations: if $f$ and $g$ are, respectively, the constants $f_{0}$ and $g_{0}$, we have the first integral $(y^{\prime})^{2}=y^{4}+f_{0}y^{2}+2g_{0}y+K$ (see also [Inc44, §14.315, p. 333]). When substituting for $y^{\prime}$ from (22), we obtain that $2y^{2}z-z^{2}+2by-2az-a^{2}$ is a first integral for (22). Its generic level surfaces are elliptic fibrations over the domain of the independent variable.
The system has the Hamiltonian
$y^{2}z+{\textstyle\frac{1}{2}z^{2}+az-by},$
equivalent to Malmquist’s Hamiltonian for the Painlevé II equation [Mal23, §33] (we recover Malmquist’s normalization in the coordinates $(\upsilon,\zeta)=(z+a,y)$). The system is polynomially equivalent to the first-order system associated to the Painlevé II equation (cf. Proposition 2).
Subcase $n=5$
The necessary condition arising from Eq. (25) at $(u,v)=(0,v_{1})$ is $b=-3a^{\prime}$. Set $a=3q$ and $b=-9q^{\prime}$. We have that w=2yw=-2y solves the equation
$w^{\prime\prime}=-ww^{\prime}+w^{3}-12qw+12q^{\prime}.$ (27)
This equation is integrated in [Pai02, p. 23, Eq. 6]: if $u$ is a solution to
$u^{\prime\prime}-6u^{2}=q^{\prime\prime}-6q^{2},$ (28)
and is different from $q$,
$w=\frac{u^{\prime}-q^{\prime}}{u-q}$ (29)
is a solution to (27). When considered as first-order equations in two variables, equations (27) and (28) are birationally equivalent, since, from (29), $u$ and $u^{\prime}$ can be expressed rationally in terms of $w$ and $w^{\prime}$:
$\begin{split}u&=\textstyle{\frac{1}{6}}w^{2}-q+\textstyle{\frac{1}{6}}w^{\prime},\\ u^{\prime}&=\textstyle{\frac{1}{6}}w^{3}-2qw+\textstyle{\frac{1}{6}}ww^{\prime}+q^{\prime},\end{split}$
@@
in which (28) becomes the first Painlevé equation or one of its degenerations. Thus, condition (30) is a necessary and sufficient condition for (27) to be free of movable critical points as well (compare with [Inc44, §14.314, pp. 331–333]). Condition (30) is also the one arising from the canonical equation (25) at $(u,v)=(0,0)$.
The system is birationally equivalent to the first-order one associated to the first Painlevé equation (cf. Proposition 2).
3.7.3. Case $B\neq 0$, n=3n=-3